OpenStudy (tiffanymak1996):

please help solve this:

OpenStudy (tiffanymak1996):

\[\lim_{n \rightarrow +\infty} \frac{ \sqrt{(n+1)} }{ \sqrt{n} }\]

OpenStudy (lgbasallote):

\[\large \implies \lim_{n \rightarrow +\infty} \sqrt{\frac{n+1}{n}}\] \[\large \implies \lim_{n \rightarrow +\infty} \sqrt{\frac nn + \frac 1n}\] \[\large \implies \lim_{n \rightarrow +\infty} \sqrt{1 + \frac 1n}\] now plug in infty \[\large \implies 1 + \frac{1}{\infty}\] \[\large \implies 1 + 0\] \[\large \implies 1\] does that help?

OpenStudy (lgbasallote):

lol i forgot to square root it...still same thing though

OpenStudy (tiffanymak1996):

thx :)

OpenStudy (lgbasallote):

should be: plug in infinity: \[\large \implies \sqrt{1 + \frac{1}{\infty}}\] \[\large \implies \sqrt{1 + 0}\] \[\large \implies \sqrt 1\] \[\large \implies 1\]

OpenStudy (lgbasallote):

welcome